# Tropical Gradient Descent

### v2

@@ Appendix C Learning Rate Tuning
The Adam algorithms — Adam, Adamax and TrAdamax — are run using the default $\beta=(0.9,0.999)$ parameters. We then tune the learning rates used for each method; Gradient Descent, Tropical Descent, SGD, Tropical SGD, Adam, Adamax, and TrAdamax. We perform an initial test of learning rates $\gamma$ on a logarithmic scale from 0.0001 to 30 over 1000 steps for each algorithm, with 5 random initializations for each dataset. From this, we narrowed our search to take 18 initializations at 0.5 intervals on the log scale in near optimal regions for each problem, gradient method, and data dimensionality. We then selected the learning rate with minimal mean log relative error in each case; Tables 3, 4, 5, 6, 7, 8 and 9 show the selected learning rates.
Table 3: Learning rates for each gradient method when solving the Fermat–Weber problem.
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 4: Learning rates for each gradient method when solving the Fréchet mean problem.
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 5: Learning rates for each gradient method when solving the 2-Wasserstein projection problem.
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 6: Learning rates for each gradient method when solving the $\infty$-Wasserstein projection problem.
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 7: Learning rates for each gradient method when solving the polytope PCA problem.
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 8: Learning rates for each gradient method when solving the logistic regression problem.
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 9: Learning rates for each gradient method when solving the linear regression problem.
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
@@ Appendix D Runtimes
Experiments were implemented on a AMD EPYC 7742 node using a single core, 8 GB. Table 10 shows the average time taken per initialization for each gradient method and optimization problem.
@@
2-Wasserstein Projection 2.739 2.134 2.392 2.209 2.341 2.246 2.272
$\infty$-Wasserstein Projection 2.947 3.049 2.772 2.565 2.501 3.162 2.807
@@ E.1 Fréchet Means
Here we include the results for experiments run for the computation of Fréchet means.
Table 11: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the Fréchet mean objective function. The minimal mean errors for each dataset are in bold.
@@
6$\times$10 Branching Data -4.59 -4.54 -3.98 -2.99 -4.58 -4.59 -4.53
6$\times$10 Coalescent Data -4.58 -4.47 -4.07 -1.82 -4.57 -4.58 -4.45
6$\times$10 Gaussian Data -4.59 -4.33 -4.09 -2.63 -4.58 -4.59 -4.35
6$\times$100 Branching Data -4.59 -4.59 -4.35 -4.03 -4.59 -4.58 -4.59
6$\times$100 Coalescent Data -4.57 -4.56 -4.36 -3.29 -4.55 -4.56 -4.57
6$\times$100 Gaussian Data -4.59 -4.60 -4.39 -4.12 -4.59 -4.58 -4.59
28$\times$10 Branching Data -4.58 -3.73 -3.78 -1.98 -4.55 -4.58 -3.83
28$\times$10 Coalescent Data -4.56 -3.67 -3.98 -2.46 -4.53 -4.57 -3.70
28$\times$10 Gaussian Data -4.55 -4.07 -3.91 -2.41 -4.55 -4.57 -4.04
28$\times$100 Branching Data -4.59 -4.58 -4.03 -2.98 -4.58 -4.59 -4.57
28$\times$100 Coalescent Data -4.54 -4.56 -4.08 -2.66 -4.47 -4.57 -4.46
28$\times$100 Gaussian Data -4.59 -4.56 -4.08 -2.78 -4.59 -4.59 -4.58
Figure 11: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the Fréchet mean objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
@@ E.2 Wasserstein Projections
From Tables 12 and 13, we see that when solving the 2-Wasserstein projection problem, tropical methods are optimal in all cases other than $28\times 10$ coalescent data. For the $\infty$-Wasserstein projection problem in small dimensions, tropical descent outperforms other methods by some margin, while the comparative performances are less consistent for the $\infty$-Wasserstein projection problem in high dimensions.
Table 12: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the $2$-Wasserstein projection objective function. The minimal mean errors for each dataset are in bold.
@@
6$\times$10 Branching Data -3.02 -4.09 -2.16 -2.11 -2.95 -3.00 -3.89
6$\times$10 Coalescent Data -3.71 -3.92 -2.44 -3.13 -3.78 -3.68 -4.12
6$\times$10 Gaussian Data -3.59 -3.83 -2.87 -2.75 -3.61 -3.57 -3.89
6$\times$100 Branching Data -3.44 -4.24 -2.91 -1.60 -3.79 -3.79 -4.02
6$\times$100 Coalescent Data -4.26 -4.55 -3.27 -2.58 -4.39 -4.45 -4.50
6$\times$100 Gaussian Data -3.69 -4.17 -3.33 -2.73 -3.78 -3.78 -4.18
28$\times$10 Branching Data -3.39 -4.31 -3.16 -2.03 -3.53 -3.55 -4.32
28$\times$10 Coalescent Data -4.11 -3.79 -3.39 -2.42 -4.17 -4.12 -3.80
28$\times$10 Gaussian Data -3.43 -3.91 -3.21 -2.15 -3.50 -3.44 -3.83
28$\times$100 Branching Data -3.61 -4.55 -3.37 -2.65 -3.85 -3.85 -4.51
28$\times$100 Coalescent Data -4.11 -4.57 -3.89 -3.32 -4.21 -4.19 -4.56
28$\times$100 Gaussian Data -3.66 -4.52 -3.41 -2.61 -3.82 -3.80 -4.49
Table 13: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the $\infty$-Wasserstein projection objective function. The minimal mean errors for each dataset are in bold.
@@
6$\times$10 Branching Data -1.63 -4.32 -0.57 -0.91 -1.68 -1.69 -4.32
6$\times$10 Coalescent Data -4.00 -4.47 -0.99 -1.43 -4.03 -4.04 -4.46
6$\times$10 Gaussian Data -3.77 -4.42 -1.76 -3.33 -3.76 -3.75 -4.40
6$\times$100 Branching Data -3.35 -4.48 -1.94 -0.95 -3.34 -3.36 -4.47
6$\times$100 Coalescent Data -3.85 -4.51 -2.06 -2.21 -3.84 -3.85 -4.51
6$\times$100 Gaussian Data -3.63 -4.40 -1.61 -1.22 -3.70 -3.63 -4.36
28$\times$10 Branching Data -3.68 -3.53 -2.22 -1.17 -4.11 -3.93 -3.17
28$\times$10 Coalescent Data -4.22 -3.95 -2.76 -1.92 -4.42 -4.36 -3.94
28$\times$10 Gaussian Data -4.02 -3.14 -2.13 -0.95 -4.43 -4.30 -2.85
28$\times$100 Branching Data -3.46 -3.58 -1.37 -1.19 -3.68 -3.59 -3.39
28$\times$100 Coalescent Data -4.41 -3.92 -2.09 -2.00 -4.54 -4.51 -3.81
28$\times$100 Gaussian Data -2.98 -3.30 -1.64 -1.61 -3.07 -3.04 -2.80

### v3

@@ Appendix C Learning Rate Tuning
The Adam algorithms—Adam, Adamax and TrAdamax—are run using the default $\beta=(0.9,0.999)$ parameters. We then tune the learning rates used for each method; CD, TD, SGD, TSGD, Adam, Adamax, and TrAdamax. We take 10 initializations at unit intervals on a log scale in the region $e^{-6},e^{4}]$ for each problem, gradient method, and data dimensionality. We then selected the learning rate with minimal mean log relative error in each case; Tables 6, 7, 8, 9, 10, 11 and 12 show the selected learning rates.
For the Auction data in Subsection 4.3.2, we take 10 initializations at unit intervals in the region of $e^{-10},e^{5}]$ for each dataset; Table 13 shows the selected learning rates which minimize the mean log relative error.
Table 6: Tuned learning rates for each gradient method when solving the Fermat–Weber problem.
Data
Dim
Data
Count
CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 7: Tuned learning rates for each gradient method when solving the Fréchet mean problem.
Data
Dim
Data
Count
CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 8: Tuned learning rates for each gradient method when solving the 2-Wasserstein projection problem.
Data
Dim
Data
Count
CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 9: Tuned learning rates for each gradient method when solving the $\infty$-Wasserstein projection problem.
Data
Dim
Data
Count
CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 10: Tuned learning rates for each gradient method when solving the polytope PCA problem.
Data
Dim
Data
Count
CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 11: Tuned learning rates for each gradient method when solving the logistic regression problem.
Data
Dim
Data
Count
CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 12: Tuned learning rates for each gradient method when solving the linear regression problem.
Data
Dim
Data
Count
CD TD SGD TSGD Adam Adamax TrAdamax
@@
Table 13: Tuned learning rates for each gradient method when solving the linear regression problem with Auction data.
Data
Dim
Data
Count
CD TD SGD TSGD Adam Adamax TrAdamax
3 6 2.72 0.368 7.39 0.0498 0.000912 0.000912 0.0183
100 0.0183 2.72 0.0183 0.0498 0.000912 0.000912 0.00248
@@ Appendix D Run Times
Experiments were implemented on a AMD EPYC 7742 node using a single core, 96 GB. Table 14 shows the average time taken per initialization for each gradient method and optimization problem.
@@
2-Wasserstein
Projection
2.739 2.134 2.392 2.209 2.341 2.246 2.272
$\infty$-Wasserstein
Projection
2.947 3.049 2.772 2.565 2.501 3.162 2.807
@@ E.1 Fréchet Means
Here we include the results for experiments run for the computation of Fréchet means. (Table 15, Figure 13)
Table 15: The mean log relative error of each gradient method after 1000 steps with tuned learning rates for branching process, coalescent process, and Gaussian datasets of size $K=10,100$ in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$, $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$ when minimizing the Fréchet mean objective function. The minimal mean errors for each dataset are in bold.
@@
[…]
Figure 13: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the Fréchet mean objective. Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
@@ E.2 Wasserstein Projections
From Tables 16 and 17, we see that when solving the 2-Wasserstein projection problem, tropical methods are optimal in all cases other than $28\times 10$ coalescent data. For the $\infty$-Wasserstein projection problem in small dimensions, tropical descent outperforms other methods by some margin, while the comparative performances are less consistent for the $\infty$-Wasserstein projection problem in high dimensions.
Table 16: The mean log relative error of each gradient method after 1000 steps with tuned learning rates for branching process, coalescent process, and Gaussian datasets of size $K=10,100$ in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$, $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$ when minimizing the $2$-Wasserstein projection objective function. The minimal mean errors for each dataset are in bold.
@@
[…]
Table 17: The mean log relative error of each gradient method after 1000 steps with tuned learning rates for branching process, coalescent process, and Gaussian datasets of size $K=10,100$ in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$, $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$ when minimizing the $\infty$-Wasserstein projection objective function. The minimal mean errors for each dataset are in bold.
@@
[…]